# OmniBal: Towards Fast Instruct-tuning for Vision-Language Models via Omniverse Computation Balance

### v3

@@ Abstract
Vision-language instruct-tuning models have recently made significant progress due to their more comprehensive understanding of the world. In this work, we discover that large-scale 3D parallel training on those models leads to an imbalanced computation load across different devices. The vision and language parts are inherently heterogeneous: their data distribution and model architecture differ significantly, which affects distributed training efficiency. To address this issue, we rebalance the computational load from data, model, and memory perspectives, achieving more balanced computation across devices. Specifically, for the data, instances are grouped into new balanced mini-batches within and across devices. A search-based method is employed for the model to achieve a more balanced partitioning. For memory optimization, we adaptively adjust the re-computation strategy for each partition to utilize the available memory fully. These three perspectives are not independent but are closely connected, forming an omniverse balanced training framework. extensive experiments are conducted to validate the effectiveness of our method. Compared with the open-source training code of InternVL-Chat, training time is reduced greatly, achieving about 1.8$\times$ speed-up. Our method’s efficacy and generalizability are further validated across various models and datasets. Codes are be released at https://github.com/ModelTC/OmniBal.
@@ 1 Introduction
Large language models (LLM) have brought new possibilities to many fields. Multi-modal models, particularly Vision-Language Models (VLMs) (Alayrac et al., 2022; Team et al., 2023a; Reid et al., 2024; Liu et al., 2023a; Bai et al., 2023b; Chen et al., 2023), are advancing rapidly due to their deeper understanding of the world. The training scale of Vision-Language Models (VLMs) continues to expand, with increasingly larger datasets incorporating more text and higher-resolution images. Compared with the LLaVA-1.5  (Liu et al., 2023a), the InternVL-Chat (Chen et al., 2024) has expanded the dataset size from 665K to 5M and increased image resolution from 336x336 to 3840x2160. At the model level, larger vision encoders are adopted. The InternVL-Chat upgrades the visual encoder from $\sim$300M ViT-L-336px (Radford et al., 2021) to $\sim$6B InternViT-448px (Chen et al., 2023).
@@
In this work, we find that for instruct-tuning large vision-language models, the heterogeneous nature of data and model structures brings new challenges to 3D parallelism training: (1) Varying input sizes of LLM and VIT cause imbalance computational loads across training iterations and devices. (2) The heterogeneity between LLM and VIT models leads to inherent differences in the computational load of their transformer blocks. Along with varying input sizes, this inevitably results in uneven computational load and computational bubbles. (3) Input size variation and computational imbalance compel us to use the most aggressive re-computation (checkpointing) (Li et al., 2014) strategy to prevent program crashes, which wastes computational resources. We refer to those issues caused by the heterogeneity in data and model structures in large vision-language models as the Computation Imbalance problem, which reduces training efficiency.
@@
Figure 1: Overview of the computation imbalanced problem and our proposed solution in Standard Vision-Language instruct-tuning framework. We consider the bottleneck issues of data, model, and memory, and propose an omniverse solution addressing these three aspects, each providing the foundation for the next.
@@
To address this problem, a simple and efficient training framework called Omniverse Balance (OmniBal) is proposed, to balance computational load across multiple devices. This framework systematically balances computation in three bottlenecks, i.e. data, model, and memory, as shown in Figure 1. OmniBal works in these three closely connected aspects. Data lays the groundwork for addressing model imbalances, while data and model form the foundation for solving memory issues. Ultimately, these three aspects collaborate to achieve balanced computation. Data: The balanced dynamic mini-batch method is proposed to group instances as new mini-batches according to text length and number of images. Specifically, an iterative algorithm based on sampling and filtering combines data of different sizes into balanced groups, ensuring stable input sizes; Model: We propose balanced model partitioning to evenly spread the computational load of LLM and VIT across devices. Using a search-based approach, we efficiently find optimal partition strategies within a small search space, enabling adaptation to different model architectures and hardware platforms. The balanced dynamic mini-batch method facilitates balanced model partitioning by ensuring input sizes are consistent in advance. Memory: A balanced adaptive re-computation method is proposed to optimize the re-computation strategy on each device, maximizing both memory utilization and training speed. We calculate the memory requirements of different models to adjust the re-computation strategy adaptively. Notably, our proposed balanced dynamic mini-batch and model partitioning ensures balanced computational loads on each device, making memory analysis feasible.
@@
Extensive experiments are performed on various open-source VLM models at different scales, reducing overall training times significantly. GPU days are reduced for InternVL-Chat-1.5 (6+20B) from 61.8 to 21.3 under the Megatron-DeepSpeed  (microsoft, 2020) backend. Scaling up to InternVL-Chat-1.5-Plus (6+34B), we consistently observe great speed-up, from 75.4 to 30.5 GPU days. We conduct thorough generalization experiments, including various datasets, hardware configurations, and multiple model combinations. Consistent and substantial improvements are observed across all experiments, demonstrating the effectiveness and versatility of our method.
@@
  • We are the first to identify and address the computational imbalance problem in large-scale VLM (Vision-Language Model) instruct-tuning training, proposing a systematic solution tailored to this challenge.
@@
  • Our method achieves great speed-up while maintaining model performance on the open-source training code. Its efficacy and generalizability are further validated across various models datasets and tasks.
@@ 2.1 Multi-Modal Large Language Model(MLLM)
Large language models, such as ChatGPT (OpenAI, 2023a), GPT-4 (OpenAI, 2023b), Llama series  (Touvron et al., 2023a, b; AI, 2024), and Gemini series (Team et al., 2023b; Reid et al., 2024), have seen significant advancements recently. They rely on large datasets for training to achieve strong performance, particularly in few-shot and zero-shot scenarios. Typically, they are built on textual data and can only accept text inputs. However, real-world scenarios often involve rich multi-modal information, e.g., images. It has driven the development of large vision language models (VLMs). Visual encoders like Vision Transformer (ViT) (Dosovitskiy et al., 2021) usually incorporate vision information. A cross-modal connector is also required to align the vision encoder outputs to the language models. LLaVA (Touvron et al., 2023a) uses the simplest MLP, BLIP series  (Li et al., 2022, 2023; Dai et al., 2024) uses Q-former, Qwen-VL-Chat (Bai et al., 2023b) uses a cross-attention module. VLMs expand large language models’ capabilities and application scenarios by instruct-tuning with text and image data. However, introducing multi-modal data and heterogeneous encoders brings challenges to the training.
@@ 2.2 Large-Scale Distributed Training
Distributed training is essential for efficiently utilizing multiple GPUs to train large language models. It is achieved through 3D parallelism (Shoeybi et al., 2019; Rajbhandari et al., 2020; microsoft, 2020): data, tensor, and pipeline parallelism. Data Parallelism splits the entire dataset into mini-batches and assigns them to multiple devices, each with a model replica. This approach maximizes the use of GPU power for large datasets. DeepSpeed Zero (Rajbhandari et al., 2020) enhances it by reducing weight redundancy. However, it can still be challenged by the memory limits of individual devices when handling huge models. Tensor Parallelism distributes a model’s weight matrices across multiple devices, enabling parallel matrix operations (Shoeybi et al., 2019) and reducing per-device memory requirements. This method accelerates computation but requires dense inter-device communication, typically restricted to single-node deployments to minimize latency. Pipeline Parallelism divides a model into segments and assigns them to different devices, creating a computation flow like a production line. This technique facilitates larger model scaling across nodes. GPipe (Huang et al., 2019) proposes micro-batching to decrease forward bubbles. PipeDream (Narayanan et al., 2019) further proposes a one-forward-one-backward (1F1B) scheme to optimize memory usage. In pipeline parallelism, uneven layer partitioning can cause significant pipeline bubbles. PipeDream (Narayanan et al., 2019) and AdaPipe (Sun et al., 2024) optimize model partitioning and re-computation strategies based on profiling and dynamic programming, respectively. However, these advancements are primarily tested in text-based models and may require adaptation for large vision language model scenarios.
3 Problem Analysis
3.1 Computation Imbalance
@@
Figure 2: The Problem of Computation Imbalance in VLM Instruct-tuning Training Pipeline. DP-0 and DP-1 represent different Data Parallel processes. T-0 and T-1 represent different training times. TIME and MEM represent forward time and cost memory in the current stage respectively. STD stands for standard deviation.
@@
In this section, we explore the unique challenges of large-scale distributed training for vision-language models, focusing on two dimensions: Inter-Stage and Intra-Stage computation imbalance. Inter-Stage means the computation imbalance of different pipeline parallel stages. Intra-Stage indicates the computation imbalance of the same stage across time and device. Figure 2 shows these two computation imbalances more intuitively. And they both include three specific levels: data, model, and memory. To quantify this problem, we used the InternVL-Chat-1.2 dataset (Chen et al., 2024) to perform profile statistics shown in Tabel 1. For the Intra-Stage, we counted the information of Stage 1 as a sample.
@@
Data Imbalance: LLMs are trained on texts using next-token prediction, allowing consistent input lengths through arbitrary text sub-strings. In contrast, VLMs handle texts and images, requiring data integrity, and preventing arbitrary truncation. The varying number of images, resolutions, and text lengths result in considerable differences in input sizes across mini-batches. From Tabel 1 and Figure 2, data imbalance occurs in Inter-Stage and Intra-Stage. To better quantify the impact of dynamic input, we define the DistRatio (introduced in Section 1) to measure the degree of data imbalance of VIT and LLM.
@@
Model Imbalance: LLMs use identical transformer modules with the same computational load. Evenly dividing these layers in pipeline parallelism distributes the load effectively. However, VLMs require additional image pre-processing, necessitating an image encoder. The structural disparity between VIT and LLM results in different computational demands. From Tabel 1 and Figure 2, the standard deviation of forward time is huge in both Inter-Stage and Intra-Stage, indicating a serious computation imbalance.
@@
Differences between VLM and LLM training: As mentioned above, the difference between VLM and LLM arises from the data composition and model structure, resulting in unique Inter-Stage and Intra-Stage challenges. Inter-Stage: Since LLM has a fixed structure, the model can be equally divided and there is no Inter-Stage imbalance for any input. Dynamic input or inconsistent text-image ratios in heterogeneous VLM will lead to an Inter-Stage imbalance problem. Intra-Stage: For the LLM-Pretrain task, the input is fixed and there is no Intra-Stage imbalance problem. Dynamic input can be converted into static input by simple packing (Kosec et al., 2021) to reduce the computation imbalance for the LLM-SFT task. However, VLM instruct-tuning training cannot rely on simple packing to ensure fixed inputs for VIT and LLM, resulting in computation imbalance problems.
3.2 Quantitative Evidence of Computation Imbalance Problem
The computation imbalance of VLM instruct-tuning training consists of two dimensions: Inter-Stage and Intra-Stage:
Inter-Stage imbalance means the computation imbalance of different pipeline parallel stages (stage 1 to stage 4 in Figure 2).
Intra-Stage imbalance indicates the computation imbalance of the same stage across time and devices. They both include three specific levels: data, model, and memory.
Data Evidence: As shown in Figure 2, at time T-0, the Vision and LLM inputs for DP-0 (Group 0 of data-parallel) and DP-1 (Group 1 of data-parallel) differ (Img=4, Text=2k vs. Img=9, Text=4k). Different inputs will bring different computational complexities. For example, the total forward time of DP-0 at time T-0 is 970ms (500 + 160 + 150 + 160), and the total forward time of DP-1 at time T-0 is 1730ms (900 + 300 + 260 + 270). DP-0 needs to wait for DP-1 to complete all training before updating parameters, which will cause DP-0’s GPU resources to wait for a long time. This is the Intra-Stage imbalance problem across devices. For DP-0, the inputs at times T-0 and T-1 on the same device also differ significantly (e.g., Img=4, Text=2k versus Img=20, Text=16k), which is an Intra-Stage imbalance problem over time. This results in a substantial variance in forward time and memory usage, creating challenges in determining a globally optimal model partition strategy.
Model Evidence For DP-0, at T-0 and T-1, the standard deviation (148ms and 402 ms) of the forward time across different stages is significant, leading to serious bubbles in pipeline parallelism and slowing down training speed. This is the Inter-Stage imbalance problem. Furthermore, the computation distribution across stages shows a huge difference between T-0 and T-1. For example, the computation distribution at T-0 is (500:160:150:160)=(0.52:0.16:0.15:0.16),while the computation distribution at T-1 is (2000:1600:1000:1100)=(0.35,0.28,0.18,0.19).The varying computation distribution makes it challenging to determine the optimal model partition strategy. This includes the Inter-Stage and Intra-Stage problems (over time).
Memory Evidence The memory consumption for Stage-1 at different time points is as follows: 50 GB at T-0, 80 GB at T-1 for DP-0, 68 GB at T-0, and 60 GB at T-1 for DP-1. To prevent program crashes, we must configure the re-computation strategy based on the highest memory usage, which is 80 GB. This approach introduces additional computational overhead, categorized as the Intra-Stage imbalance problem (over time). Additionally, for DP-1, the memory usage across different stages at both T-0 and T-1 has a high standard deviation (6.5 GB and 8.4 GB, respectively). Using the same re-computation strategy for different stages will also bring computational overhead according to the highest memory usage. This variation represents the Inter-Stage imbalance problem.

### v4

# OmniBal: Towards Fast Instruction-Tuning for Vision-Language Models via Omniverse Computation Balance

@@ Abstract
Vision-language instruction-tuning models have recently achieved significant performance improvements. In this work, we discover that large-scale 3D parallel training on those models leads to an imbalanced computation load across different devices. The vision and language parts are inherently heterogeneous: their data distribution and model architecture differ significantly, which affects distributed training efficiency. To address this issue, we rebalance the computational load from data, model, and memory perspectives, achieving more balanced computation across devices. Specifically, for the data, instances are grouped into new balanced mini-batches within and across devices. A search-based method is employed for the model to achieve a more balanced partitioning. For memory optimization, we adaptively adjust the re-computation strategy for each partition to utilize the available memory fully. These three perspectives are not independent but are closely connected, forming an omniverse balanced training framework. Extensive experiments are conducted to validate the effectiveness of our method. Compared with the open-source training code of InternVL-Chat, training time is reduced greatly, achieving about 1.8$\times$ speed-up. Our method’s efficacy and generalizability are further validated across various models and datasets. Codes will be released at https://github.com/ModelTC/OmniBal.
@@ 1 Introduction
Large language models (LLMs) have brought new possibilities to many fields. Multi-modal models, particularly Vision-Language Models (VLMs) (Alayrac et al., 2022; Team et al., 2023a; Reid et al., 2024; Liu et al., 2023a; Bai et al., 2023b; Chen et al., 2023), are advancing rapidly due to their deeper understanding of the world. The training scale of Vision-Language Models (VLMs) continues to expand, with increasingly larger datasets incorporating more text and higher-resolution images. Compared with the LLaVA-1.5 (Liu et al., 2023a), the InternVL-Chat (Chen et al., 2024) has expanded the dataset size from 665K to 5M and increased image resolution from 336x336 to 3840x2160. At the model level, larger vision encoders are adopted. The InternVL-Chat upgrades the visual encoder from $\sim$300M ViT-L-336px (Radford et al., 2021) to $\sim$6B InternViT-448px (Chen et al., 2023).
@@
In this work, we find that for instruction-tuning large vision-language models, the heterogeneous nature of data and model structures brings new challenges to 3D parallelism training: (1) Varying input sizes of LLM and VIT cause imbalanced computational loads across training iterations and devices. (2) The heterogeneity between LLM and VIT models leads to inherent differences in the computational load of their transformer blocks. Along with varying input sizes, this inevitably results in uneven computational load and computational bubbles. (3) Input size variation and computational imbalance compel us to use the most aggressive re-computation (checkpointing) (Li et al., 2014) strategy to prevent program crashes, which wastes computational resources. We refer to those issues caused by the heterogeneity in data and model structures in large vision-language models as the Computation Imbalance problem, which reduces training efficiency.
@@
Figure 1: Overview of the computation imbalanced problem and our proposed solution in the Standard Vision-Language instruction-tuning framework. We consider the bottleneck issues of data, model, and memory, and propose an omniverse solution addressing these three aspects, each providing the foundation for the next.
@@
To address this problem, a simple and efficient training framework called Omniverse Balance (OmniBal) is proposed to balance computational load across multiple devices. This framework systematically balances computation in three bottlenecks, i.e. data, model, and memory, as shown in Figure 1. OmniBal works in these three closely connected aspects. Data lays the groundwork for addressing model imbalances, while data and model form the foundation for solving memory issues. Ultimately, these three aspects collaborate to achieve balanced computation. Data: The balanced dynamic mini-batch method is proposed to group instances as new mini-batches according to text length and number of images. Specifically, an iterative algorithm based on sampling and filtering combines data of different sizes into balanced groups, ensuring stable input sizes; Model: We propose balanced model partitioning to evenly spread the computational load of LLM and VIT across devices. Using a search-based approach, we efficiently find optimal partition strategies within a small search space, enabling adaptation to different model architectures and hardware platforms. The balanced dynamic mini-batch method facilitates balanced model partitioning by ensuring input sizes are consistent in advance. Memory: A balanced adaptive re-computation method is proposed to optimize the re-computation strategy on each device, maximizing both memory utilization and training speed. We calculate the memory requirements of different models to adjust the re-computation strategy adaptively. Notably, our proposed balanced dynamic mini-batch and model partitioning ensures balanced computational loads on each device, making memory analysis feasible.
@@
Extensive experiments are performed on various open-source VLM models at different scales, reducing overall training times significantly. GPU days are reduced for InternVL-Chat-1.5 (6+20B) from 61.8 to 21.3 under the Megatron-DeepSpeed  (microsoft, 2020) backend. Scaling up to InternVL-Chat-1.5-Plus (6+34B), we consistently observe a great speed-up, from 75.4 to 30.5 GPU days. We conduct thorough generalization experiments, including various datasets, hardware configurations, and multiple model combinations. Consistent and substantial improvements are observed across all experiments, demonstrating the effectiveness and versatility of our method.
@@
  • We are the first to identify and address the computational imbalance problem in large-scale VLM (Vision-Language Model) instruction-tuning training, proposing a systematic solution tailored to this challenge.
@@
  • Our method achieves a great speed-up while maintaining model performance on the open-source training code. Its efficacy and generalizability are further validated across various models datasets and tasks.
@@ 2.1 Multi-Modal Large Language Model(MLLM)
Large language models, such as ChatGPT (OpenAI, 2023a), GPT-4 (OpenAI, 2023b), Llama series  (Touvron et al., 2023a, b; AI, 2024), and Gemini series (Team et al., 2023b; Reid et al., 2024), have seen significant advancements recently. They rely on large datasets for training to achieve strong performance, particularly in few-shot and zero-shot scenarios. Typically, they are built on textual data and can only accept text inputs. However, real-world scenarios often involve rich multi-modal information, e.g., images. It has driven the development of large vision language models (VLMs). Visual encoders like Vision Transformer (ViT) (Dosovitskiy et al., 2021) usually incorporate vision information. A cross-modal connector is also required to align the vision encoder outputs to the language models. LLaVA (Touvron et al., 2023a) uses the simplest MLP, BLIP series  (Li et al., 2022, 2023; Dai et al., 2024) uses Q-former, Qwen-VL-Chat (Bai et al., 2023b) uses a cross-attention module. VLMs expand large language models’ capabilities and application scenarios by instruction-tuning with text and image data. However, introducing multi-modal data and heterogeneous encoders brings challenges to the training.
@@ 2.2 Large-Scale Distributed Training
Distributed training is essential for efficiently utilizing multiple GPUs to train large language models. It is achieved through 3D parallelism (Shoeybi et al., 2019; Rajbhandari et al., 2020; microsoft, 2020): data, tensor, and pipeline parallelism. Data Parallelism splits the entire dataset into mini-batches and assigns them to multiple devices, each with a model replica. This approach maximizes the use of GPU power for large datasets. DeepSpeed Zero (Rajbhandari et al., 2020) enhances it by reducing weight redundancy. However, it can still be challenged by the memory limits of individual devices when handling huge models. Tensor Parallelism distributes a model’s weight matrices across multiple devices, enabling parallel matrix operations (Shoeybi et al., 2019) and reducing per-device memory requirements. This method accelerates computation but requires dense inter-device communication, typically restricted to single-node deployments to minimize latency. Pipeline Parallelism divides a model into segments and assigns them to different devices, creating a computation flow like a production line. This technique facilitates larger model scaling across nodes. GPipe (Huang et al., 2019) proposes micro-batching to decrease forward bubbles. PipeDream (Narayanan et al., 2019) further proposes a one-forward-one-backward (1F1B) scheme to optimize memory usage. In pipeline parallelism, uneven layer partitioning can cause significant pipeline bubbles. PipeDream (Narayanan et al., 2019) and AdaPipe (Sun et al., 2024) optimize model partitioning and re-computation strategies based on profiling and dynamic programming, respectively. However, these advancements are primarily tested in text-based models and may require adaptation for large vision-language model scenarios.
3 Computation Imbalance
@@
Figure 2: The Problem of Computation Imbalance in VLM Instruction-Tuning Training Pipeline. DP-0 and DP-1 represent different Data Parallel processes. T-0 and T-1 represent different training times. TIME and MEM represent forward time and cost memory in the current stage, respectively. STD stands for standard deviation.
@@
In this section, we explore the unique challenges of large-scale distributed training for vision-language models, focusing on two dimensions: Inter-Stage and Intra-Stage computation imbalance. Inter-Stage means the computation imbalance of different pipeline parallel stages. Intra-Stage indicates the computation imbalance of the same stage across time and devices. Figure 2 shows these two computation imbalances more intuitively. And they both include three specific levels: data, model, and memory. To quantify this problem, we used the InternVL-Chat-1.2 dataset (Chen et al., 2024) to perform profile statistics shown in Table 1. For the Intra-Stage, we counted the information of Stage 1 as a sample.
@@
Data Imbalance: LLMs are trained on texts using next-token prediction, allowing consistent input lengths through arbitrary text sub-strings. In contrast, VLMs handle texts and images, requiring data integrity, and preventing arbitrary truncation. The varying number of images, resolutions, and text lengths result in considerable differences in input sizes across mini-batches. From Table 1 and Figure 2, data imbalance occurs in Inter-Stage and Intra-Stage. To better quantify the impact of dynamic input, we define the DistRatio (introduced in Section 1) to measure the degree of data imbalance of VIT and LLM.
Data Imbalance Evidence: As shown in Figure 2, at time T-0, the Vision and LLM inputs for DP-0 (Group 0 of data-parallel) and DP-1 (Group 1 of data-parallel) differ (Img=4, Text=2k vs. Img=9, Text=4k). Different inputs will bring different computational complexities. For example, the total forward time of DP-0 at time T-0 is 970ms (500 + 160 + 150 + 160), and the total forward time of DP-1 at time T-0 is 1730ms (900 + 300 + 260 + 270). DP-0 needs to wait for DP-1 to complete all training before updating parameters, which will cause DP-0’s GPU resources to wait for a long time. This is the Intra-Stage imbalance problem across devices. For DP-0, the inputs at times T-0 and T-1 on the same device also differ significantly (e.g., Img=4, Text=2k versus Img=20, Text=16k), which is an Intra-Stage imbalance problem over time. This results in a substantial variance in forward time and memory usage, creating challenges in determining a globally optimal model partition strategy.
@@
Model Imbalance: LLMs use identical transformer modules with the same computational load. Evenly dividing these layers in pipeline parallelism distributes the load effectively. However, VLMs require additional image pre-processing, necessitating an image encoder. The structural disparity between VIT and LLM results in different computational demands. From Table 1 and Figure 2, the standard deviation of forward time is huge in both Inter-Stage and Intra-Stage, indicating a serious computation imbalance.
Model Imbalance Evidence: As shown in Figure 2, for DP-0, at T-0 and T-1, the standard deviation (148ms and 402 ms) of the forward time across different stages is significant, leading to serious bubbles in pipeline parallelism and slowing down training speed. This is the Inter-Stage imbalance problem. Furthermore, the computation distribution across stages shows a huge difference between T-0 and T-1. For example, the computation distribution at T-0 is (500:160:150:160)=(0.52:0.16:0.15:0.16),while the computation distribution at T-1 is (2000:1600:1000:1100)=(0.35,0.28,0.18,0.19).The varying computation distribution makes determining the optimal model partition strategy challenging. This includes the Inter-Stage and Intra-Stage problems (over time).
@@
Memory Imbalance Evidence: As shown in Figure 2, the memory consumption for Stage-1 at different time points is as follows: 50 GB at T-0, 80 GB at T-1 for DP-0, 68 GB at T-0, and 60 GB at T-1 for DP-1. To prevent program crashes, we must configure the re-computation strategy based on the highest memory usage, which is 80 GB. This approach introduces additional computational overhead, categorized as the Intra-Stage imbalance problem (over time). Additionally, for DP-1, the memory usage across different stages at both T-0 and T-1 has a high standard deviation (6.5 GB and 8.4 GB, respectively). Using the same re-computation strategy for different stages will also bring computational overhead, according to the highest memory usage. This variation represents the Inter-Stage imbalance problem.
Differences between VLM and LLM training: As mentioned above, the difference between VLM and LLM arises from the data composition and model structure, resulting in unique Inter-Stage and Intra-Stage challenges. Inter-Stage: Since LLM has a fixed structure, the model can be equally divided, and there is no Inter-Stage imbalance for any input. Dynamic input or inconsistent text-image ratios in heterogeneous VLM will lead to an Inter-Stage imbalance problem. Intra-Stage: For the LLM-Pretrain task, the input is fixed, and there is no Intra-Stage imbalance problem. Dynamic input can be converted into static input by simple packing (Kosec et al., 2021) to reduce the computation imbalance for the LLM-SFT task. However, VLM instruction-tuning training cannot rely on simple packing to ensure fixed inputs for VIT and LLM, resulting in computation imbalance problems.